The perimeter of a parallelogram is 50 inches. If the length of one side is 15 ​inches, what is the length of a side adjacent to​ it?
Answer by Timnewman(323)

The perimeter =2(L+w)
since p=50,L=15
2(L+w)=50
2(15+w)=50
30+2w=50
2w=50-30
2w=20
w=10
the adjecent side is 10 inches.
---TIMNEWMAN-----